OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±4) and asymptotes at y = ±1/4x.

OpenStudy (anonymous):

Jdoe! Thanks lol

OpenStudy (jdoe0001):

notice, your vertices are at (0, 4) and (0, -4) the "x" doesn't change, the "y" coordinate does that means the hyperbola is moving upward/downward

OpenStudy (jdoe0001):

the distance between one vertex to the other is 4 units up plus 4 units down, so 8 units, so half-way in between, is the center, from (0, 4) to (0, -4), puts you in the origin moving upwards/downwads by 4 units to the vertices, means, THAT IS the TRAVERSE AXIS, and thus the "a" component of the hyperbola so a = 4 center = (0, 0)

OpenStudy (jdoe0001):

since the traverse axis is the "y" axis, that means the fraction with the POSITIVE sign will be the fraction with the "y", so thus far it'd look like $$ \cfrac{(y-0)^2}{4}-\cfrac{(x-0)^2}{b} = 1 $$

OpenStudy (anonymous):

Hmmm okay i see

OpenStudy (anonymous):

y^2/16 - x^2/64 = 1

OpenStudy (jdoe0001):

a = 4 \(a^2 = 2^2\)

OpenStudy (anonymous):

so the 16 is correct

OpenStudy (jdoe0001):

wait.. that made no sense lol

OpenStudy (anonymous):

lololololl

OpenStudy (jdoe0001):

right it'd be 16 hehe

OpenStudy (anonymous):

lol so what do i dooo? if a = 4 then a^2 = 16, meaning that the 16 is right

OpenStudy (anonymous):

is the 64 right?

OpenStudy (anonymous):

or should it be 256?:

OpenStudy (jdoe0001):

well, dunno thus far, you see the asymptotes for a vertical hyperbola are at \(y = k \pm \cfrac{a}{b}(x-h)\\\)

OpenStudy (anonymous):

:l i think im just gonna guess... been on this one Q for 35 mins

OpenStudy (jdoe0001):

which leads me to think b is 1

OpenStudy (jdoe0001):

if he slope of the asymptote is 1/4, and a = 4, then b has to be the numerator